Problems with DNS have now been resolved. We believe the system is now running normally.
The facility housing our East datacentre experienced DNS connectivity problems on outbound connections starting at around 16.30 UTC which gradually worsened. We switched East traffic to our South datacentre where DNS connectivity is fine. We have a team on site at the East facility determining the reason for degradation of DNS, and we are monitoring our other datacentres.
Please contact Bango if you experience any abnormal behavior